Accelerator Update for January 26, 2009 - January 28, 2009
Monday Jan 26
The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (371.8E10), and with the Antiproton Source stacking (38.9E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting on the next store.
At 8:34 AM, Operations put the shot setup on hold. The D0 experiment needed a two-hour access to repair a power supply that was limiting their ability to take data.
Operations prepared D0 for access at 8:38 AM.
Operations prepared the Pelletron for a 15-minute access at 10:03 AM.
D0 came out of access at 10:05 AM.
Operations put the TeV through a dry squeeze at 10:20 AM.
Operations put the TeV through a wet squeeze at 11:02 AM.
Operations put the TeV into shot setup at noon.
At 12:35 PM, a Linac quadrupole power supply’s (QPS513) output current was drifting. It was adjusted.
Operations halted beam to MI, Pbar, and NuMI at 1:10 PM to allow the following work: MIRF8 fan repair, MI-20 Lower power supply taken out of circuit, MI kicker (I:KPS62A) investigation, and MI-20 LCW pump work.
Operations established store 6755 at 1:26 PM with an initial luminosity of 318.55E30.
At 1:40 PM, Operations reported finding the Tevatron Electron Lens (TEL2) off. Operators contacted an expert and then turned the Lens back on.
At 2:29 PM, the MI Low Level RF system went down. An expert investigated.
Operations reestablished beam to Pbar and NuMI at 2:52 PM.
At 3 PM, MIRF18 tripped off. It reset.
A Pbar expert conducted a 20-minute study at 3:39 PM.
Operations reported at 4:38 PM that the Recycler had stashed antiprotons once during the day shift.
At 9:05 PM, LRF4 tripped off. It reset.
At 10:04 PM, Operations reported a damper trip in between transfers of antiprotons to the Recycler. The Recycler lost ~120E10 of antiprotons from the stash. An expert investigated, but he could not find anything wrong.
Operations reported at 11:59 PM that the Recycler had stashed six times during the evening shift.
Tuesday Jan 27
The midnight shift began with Operations monitoring store 6755 (95.4E30) and the stash (175.5E10), with Pbar stacking (27.7E10), and with NuMI and MiniBooNE taking beam.
At 1:18 AM, Operations lowered the NuMI beam intensity due to MI tune problems.
At 4:57 AM, the Recycler suffered another damper trip in between transfers of antiprotons. Operations contacted experts. The Recycler lost 130E10 of antiprotons.
The Run Coordinator told Operations at 5:56 AM to hold on to this store for a bit longer than normal so that Recycler experts could continue searching for the cause of their damper problems.
A Recycler expert reported at 7:21 AM that he had found a phase error in their damper system.
Operations reported at 8:01 AM that the Recycler had stashed five times during the midnight shift.
Operations terminated store 6755 at 9:13 AM.
The Neutron Therapy Facility reported at 10:06 AM that they had finished treating their last patient for the day.
Operations put the TeV into shot setup at 10:15 AM.
At 11:31 AM, Operations halted MI beam to start the following work: bringing the MI-20 Lower Power Supply back into the circuit, making up water for the MI RF and magnet LCW systems, Accumulator kicker (A:IKIK) remote control work, and installing the Booster Low Level RF ramp card upgrade.
Operations established store 6756 at 11:47 AM with an initial luminosity of 239.0E30.
MI and Pbar experts reported that all of their work was completed. Operations resumed sending beam to NuMI and Pbar at noon.
At 12:05 PM, Operations reported that the MI beam permit was tripping off due to problems with a power supply’s (E:V100) regulation. An expert investigated.
The MI power supply expert did not find a problem with E:V100. Operations resumed sending beam to NuMI and Pbar at 1:04 PM.
Operations reported at 4:03 PM that the Recycler had stashed four times during the day shift.
Recycler experts completed their damper work at 5 PM.
At 8:47 PM, the P1 transfer line beam permit tripped off for the first of two times in about an hour due to a current drift on a power supply (I:V714). It reset.
At 11:03 PM, Operations reported that the MI beam was unstable. Operators investigated and discovered a recent horizontal orbit smooth was bad. They reverted back to the settings before the smooth.
Operations reported at midnight that the Recycler had stashed eight times during the evening shift.
Wednesday Jan 28
The midnight shift began with Operations monitoring store 6756 (62.25E30) and the stash (312.5E10), with Pbar stacking (28.9E10), and with NuMI and MiniBooNE taking beam.
Operations terminated store 6756 at 3:43 AM.
Operations put the TeV into shot setup at 3:54 AM.
Operations established store 6757 at 5:24 AM with an initial luminosity of 322.65E30.
Operations halted the NuMI beam at 5:25 AM and prepared MI-65 for a scheduled access.
At 5:27 AM, MIRF12 began tripping off on gap envelope faults. It would reset, but would not stay on. Operators bypassed the station.
At 5:38 AM, Operations reported that TEL2 did not come at the beginning of the store. Operators managed to get it running properly.
Operations reported at 7:53 AM that the Recycler had stashed seven times during the midnight shift.
The Plans for Wednesday
The plans for today are to stack, stash, and store. The Run Coordinator announced that there would be a two-hour access on Thursday morning for Booster work. Some of the other machines will be allowed to do maintenance during this short shutdown.
Comments and Suggestions
|last modified 01/28/2009 email Fermilab|
|Security, Privacy, Legal|